# On finite generating sets of infinitely generated ideals

### v1

@@ 1. Introduction
The Hilbert basis theorem [Hilbert] asserts that any polynomial ideal is finitely generated. Nonetheless, determining a finite generating set for an ideal derived from infinitely many polynomials poses significant theoretical and computational challenges. For example, take the ideal of the polynomial ring ${\mathbb{Q}}[x,y,z,w]$ generated by infinitely many polynomials
@@
Later, we will demonstrate that this ideal is finitely generated by $f_{0},f_{1},f_{2},f_{3}$ (Example 3.2).
@@
Then $I$ is generated by $x^{M}$ where $M=\min\{\varphi(n_{1},\dots,n_{p})^{2}\mid n_{1},\dots,n_{p}\in{\mathbb{Z}}_{\geq 0}\}$. In particular, $I={\mathbb{Q}}[x]$ if and only if the Diophantine equation $\varphi(n_{1},\dots,n_{p})=0$ has a non-negative integer solution. Therefore, determining a generating set of $I$ depends on the solvability of the Diophantine equation $\varphi(n_{1},\dots,n_{p})=0$. According to the MRDP (Matiyasevich-Robinson-Davis-Putnam) theorem [Matiyasevich], no algorithm exists to decide this in general.
@@
Fix positive integers $p,r$, and $N$. Let $R$ be a commutative ring, and $x_{1},\dots,x_{r}\in R$. Let
@@
$\ell_{i}=\sum_{k=1}^{N}(\mathop{\mathrm{deg}}\nolimits_{n_{i}}c_{k}+1)-1.$
Then $I$ is generated by $\{f_{n_{1},\dots,n_{p}}\mid\forall i,0\leq n_{i}\leq\ell_{i}\}$.
@@ 2. Main theorem
Let $R$ be a commutative ring (not necessarily Noetherian). For $\bm{x}=(x_{1},\dots,x_{r})\in R^{r}$ and an exponent vector $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{r})\in{\mathbb{Z}}_{\geq 0}^{r}$, we use the multi-index notation $\bm{x^{\alpha}}=x_{1}^{\alpha_{1}}\cdots x_{r}^{\alpha_{r}}$. For example, let $R={\mathbb{Q}}[x,y]$, $\bm{z}=(2,x,x+y)\in R^{3}$, $\bm{\alpha}=(n+2,n,2n+1)$, and $\bm{\beta}=(n,0,0)$, then $\bm{z}^{\bm{\alpha}}+\bm{z}^{\bm{\beta}}=2^{n+2}x^{n}(x+y)^{2n+1}+2^{n}$.
@@
Let ${\mathbb{Q}}[w]$ be a polynomial ring, and $k\in{\mathbb{Z}}_{\geq 0}$.
$\binom{w}{k}=\frac{w(w-1)\dots(w-k+1)}{k!}\in{\mathbb{Q}}[w]$
is called the binomial coefficient polynomial.
For $n,k\in{\mathbb{Z}}_{\geq 0}$, the binomial coefficient $\binom{n}{k}$ represents the number of ways to choose $k$ elements from a set of $n$ elements without regarding the order of selection, and thus $\binom{n}{k}\in{\mathbb{Z}}_{\geq 0}$. It is well-known that the generating functions of binomial coefficients are rational functions.
@@
It suffices to prove the case where $R={\mathbb{Z}}$.
$\sum_{n=0}^{\infty}\binom{n}{k}t^{n}$ $=\dfrac{t^{k}}{k!}\sum_{n=0}^{\infty}n\cdot(n-1)\cdot(n-k+1)t^{n-k}$
$=\dfrac{t^{k}}{k!}\dfrac{d^{k}}{dt^{k}}\sum_{n=0}^{\infty}t^{n}=\dfrac{t^{k}}{k!}\dfrac{d^{k}}{dt^{k}}\dfrac{1}{1-t}=\dfrac{t^{k}}{(1-t)^{k+1}}.$
@@
Fix positive integers $p,r$, and $N$. Let $R$ be a commutative ring, and $\bm{x}=(x_{1},\dots,x_{r})\in R^{r}$. Let $c_{k}(\bm{n})\in R[\bm{n}]=R[n_{1},\dots,n_{p}]$ be polynomials in the parameters $\bm{n}=(n_{1},\dots,n_{p})$, and $\bm{A}^{(k)}=(\alpha^{(k)}_{ij})\in{\mathbb{Z}}_{\geq 0}^{r\times p},\bm{\beta}^{(k)}=(\beta^{(k)}_{i})\in{\mathbb{Z}}^{r}_{\geq 0}$ for $1\leq k\leq N$. Let
$f_{\bm{n}}=\sum_{k=0}^{N}c_{k}(\bm{n})\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}},~{}~{}\ell_{i}=\sum_{k=1}^{N}(\mathop{\mathrm{deg}}\nolimits_{n_{i}}c_{k}+1)-1,$
and $\bm{\ell}=(\ell_{1},\dots,\ell_{p})$.
@@
For each $k$, the coefficient polynomial $c_{k}(\bm{n})$ can be written as an $R$-linear summation of binomial coefficient polynomials:
$c_{k}(\bm{n})=\sum_{\begin{subarray}{c}\bm{m}=({m}_{1},\dots,{m}_{p})\in{\mathbb{Z}}_{\geq 0}^{p}\\ \forall j,0\leq{m}_{j}\leq\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}\end{subarray}}\gamma_{\bm{m}}\prod_{j=1}^{p}\binom{n_{j}}{{m}_{j}}$
@@
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}\prod_{j=1}^{p}\binom{n_{j}}{{m}_{j}}\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}}\bm{t}^{\bm{n}}$ $=\bm{x}^{\bm{\beta}^{(k)}}\prod_{j=1}^{p}\left\{\sum_{n_{j}=0}^{\infty}\binom{n_{j}}{{m}_{j}}\left(\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{n_{j}}\right\}$
$=\frac{\bm{x}^{\bm{\beta}^{(k)}}\bm{t}^{\bm{m}}}{\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{{m}_{j}+1}}$
@@
$=\sum_{k=1}^{N}\sum_{\begin{subarray}{c}\bm{m}=({m}_{1},\dots,{m}_{p})\in{\mathbb{Z}}_{\geq 0}^{p}\\ \forall j,0\leq{m}_{j}\leq\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}\end{subarray}}\frac{\gamma_{\bm{m}}\bm{x}^{\bm{\beta}^{(k)}}\bm{t}^{\bm{m}}}{\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{{m}_{j}+1}}.$
@@
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}f_{\bm{n}}\bm{t}^{\bm{n}}=\frac{g(\bm{t})}{\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1}},$
@@
$\mathop{\mathrm{deg}}\nolimits_{t_{j}}g(\bm{t})<\mathop{\mathrm{deg}}\nolimits_{t_{j}}\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1}=\sum_{j=1}^{p}\left(\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1\right).$
This shows that any term $\bm{t}^{\bm{m}}$ appearing in $g(\bm{t})$ with non-zero coefficient satisfies $\bm{m}\leq\bm{\ell}$. Since the denominator $\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1}$ is invertible in $R\llbracket\bm{t}\rrbracket$, the ideal $\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}^{r}_{\geq 0}\rangle\subset R$ is generated by $\{f_{\bm{n}}\mid\bm{n}\leq\bm{\ell}\}$ by Lemma 2.1. ∎
@@
for $n\in{\mathbb{Z}}_{\geq 0}$. Since degree of the coefficients $1,1,1,1$ are all 00, we have $\ell=4-1=3$. By Theorem 2.5, $\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $f_{0},\dots,f_{3}$.
@@ 3. The case of higher degree exponents
and $I=\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}\rangle$. We consider a finite generating set of $I$. Fix a sutable $\bm{k}_{0}\in{\mathbb{Z}}_{\geq 0}^{r}$ and let $J=\langle f_{\bm{n}}\mid\bm{n}\leq\bm{k}_{0}\rangle$. Then take a sufficient large $\bm{k}_{1}\in{\mathbb{Z}}_{\geq 0}^{r}$ with $\bm{k}_{0}<\bm{k}_{1}$, and compute the list

### v2

@@ Abstract.
This work has been supported by JSPS Grant-in-Aid for Scientific Research (C) (No. 22K03334).
@@ 1. Introduction
The Hilbert basis theorem [Hilbert] asserts that any polynomial ideal is finitely generated. Nonetheless, determining a finite generating set for an ideal derived from infinitely many polynomials poses significant theoretical and computational challenges. In their work [BE], Buchberger and Elías demonstrate that the ideal generated by the Fermat polynomials $f_{n}=x^{n}+y^{n}-z^{n}$ for $n\geq 3$ is, in fact, generated by $f_{3}$, $f_{4}$, and $f_{5}$. In this paper, we discuss more general situations. For example, take the ideal of the polynomial ring ${\mathbb{Q}}[x,y,z,w]$ generated by infinitely many polynomials
@@
Later, we will show that this ideal is finitely generated by $f_{0},f_{1},f_{2},$ and $f_{3}$ (Example 3.2).
@@
Then II is generated by $x^{M}$ where $M=\min\{\varphi(n_{1},\dots,n_{p})^{2}\mid n_{1},\dots,n_{p}\in{\mathbb{Z}}_{\geq 0}\}$. In particular, $I={\mathbb{Q}}[x]$ if and only if the Diophantine equation $\varphi(n_{1},\dots,n_{p})=0$ has a non-negative integer solution. Therefore, determining a generating set of II depends on the solvability of the Diophantine equation $\varphi(n_{1},\dots,n_{p})=0$. According to the MRDP (Matiyasevich-Robinson-Davis-Putnam) theorem [Matiyasevich], no algorithm exists to decide this in general.
@@
Fix positive integers $p,r$, and NN. Let $R$ be a commutative ring, and $x_{1},\dots,x_{r}\in R$. Let
@@
$\ell_{i}=\sum_{k=1}^{N}(\deg_{n_{i}}c_{k}+1)-1.$
Then II is generated by $\{f_{n_{1},\dots,n_{p}}\mid\forall i,0\leq n_{i}\leq\ell_{i}\}$.
@@ 2. Main theorem
Let $R$ be a commutative ring (not necessarily Noetherian). For $\bm{x}=(x_{1},\dots,x_{r})\in R^{r}$ and an exponent vector $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{r})\in{\mathbb{Z}}_{\geq 0}^{r}$, we use the multi-index notation $\bm{x^{\alpha}}=x_{1}^{\alpha_{1}}\cdots x_{r}^{\alpha_{r}}$, and $\bm{\alpha!}=(\alpha_{1}!)\times\dots\times(\alpha_{r}!)$. For example, let $R={\mathbb{Q}}[x,y]$, $\bm{z}=(2,x,x+y)\in R^{3}$, $\bm{\alpha}=(n+2,n,2n+1)$, and $\bm{\beta}=(n,0,0)$, then $\bm{z}^{\bm{\alpha}}+\bm{z}^{\bm{\beta}}=2^{n+2}x^{n}(x+y)^{2n+1}+2^{n}$.
@@
For a variable $w$ and $k\in{\mathbb{Z}}_{\geq 0}$, we define
$[w]_{k}=w(w-1)\dots(w-k+1).$
It is well-known that the generating function of $n]_{k},~n\in{\mathbb{Z}}_{\geq 0},$ is a rational function.
@@
It suffices to prove the case where $R={\mathbb{Z}}$. By applying $t^{k}\dfrac{d^{k}}{dt^{k}}$ on the both sides of
$\sum_{n=0}^{\infty}t^{n}=\dfrac{1}{1-t},$
we conclude the assertion. ∎
The characteristic of $R$, denoted by ${\mathrm{char}}R$, is the non-negative generator of the kernel of the unique ring homomorphism ${\mathbb{Z}}\to R$. When ${\mathrm{char}}R>0$ and $k\geq{\mathrm{char}}R$, it holds that $k!=0$ and $n]_{k}=0$ for all $n\in{\mathbb{Z}}_{\geq 0}$ in $R$. Thus, in this case, $\sum_{n=0}^{\infty}[n]_{k}t^{n}=\frac{k!t^{k}}{(1-t)^{k+1}}=0$.
@@
Fix positive integers $p,r$, and NN. Let $R$ be a commutative ring, and $\bm{x}=(x_{1},\dots,x_{r})\in R^{r}$. Let $c_{k}(\bm{n})\in R[\bm{n}]=R[n_{1},\dots,n_{p}]$ be polynomials in the parameters $\bm{n}=(n_{1},\dots,n_{p})$, and $\bm{A}^{(k)}=(\alpha^{(k)}_{ij})\in{\mathbb{Z}}_{\geq 0}^{r\times p},\bm{\beta}^{(k)}=(\beta^{(k)}_{i})\in{\mathbb{Z}}^{r}_{\geq 0}$ for $1\leq k\leq N$. Let
$f_{\bm{n}}=\sum_{k=0}^{N}c_{k}(\bm{n})\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}},~~\ell_{j}=\sum_{k=1}^{N}(\deg_{n_{j}}c_{k}+1)-1,$
and $\bm{\ell}=(\ell_{1},\dots,\ell_{p})$ where $\deg_{n_{j}}c_{k}$ denotes the degree of $c_{k}$ in $n_{j}$.
@@
For each $k$, the coefficient polynomial $c_{k}(\bm{n})$ can be written as
$c_{k}(\bm{n})=c_{k}(n_{1},\dots,n_{p})=\sum_{\begin{subarray}{c}\bm{m}=({m}_{1},\dots,{m}_{p})\in{\mathbb{Z}}_{\geq 0}^{p}\\ \forall j,0\leq{m}_{j}\leq\deg_{n_{j}}c_{k}\end{subarray}}\gamma_{\bm{m}}\prod_{j=1}^{p}[n_{j}]_{{m}_{j}}$
@@
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}\prod_{j=1}^{p}[n_{j}]_{{m}_{j}}\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}}\bm{t}^{\bm{n}}$ $=\bm{x}^{\bm{\beta}^{(k)}}\prod_{j=1}^{p}\left\{\sum_{n_{j}=0}^{\infty}[n_{j}]_{{m}_{j}}\left(\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{n_{j}}\right\}$
$=\frac{\bm{m!}\bm{x}^{\bm{\beta}^{(k)}}\bm{t}^{\bm{m}}}{\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{{m}_{j}+1}}$
@@
$=\sum_{k=1}^{N}\sum_{\begin{subarray}{c}\bm{m}=({m}_{1},\dots,{m}_{p})\in{\mathbb{Z}}_{\geq 0}^{p}\\ \forall j,0\leq{m}_{j}\leq\deg_{n_{j}}c_{k}\end{subarray}}\frac{\bm{m!}\gamma_{\bm{m}}\bm{x}^{\bm{\beta}^{(k)}}\bm{t}^{\bm{m}}}{\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{{m}_{j}+1}}.$
@@
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}f_{\bm{n}}\bm{t}^{\bm{n}}=\frac{g(\bm{t})}{\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\deg_{n_{j}}c_{k}+1}},$
@@
$\deg_{t_{j}}g(\bm{t})<\deg_{t_{j}}\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\deg_{n_{j}}c_{k}+1}=\sum_{j=1}^{p}\left(\deg_{n_{j}}c_{k}+1\right).$
This shows that any term $\bm{t}^{\bm{m}}$ appearing in $g(\bm{t})$ with non-zero coefficient satisfies $\bm{m}\leq\bm{\ell}$. Since the denominator $\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\deg_{n_{j}}c_{k}+1}$ is invertible in $R\llbracket\bm{t}\rrbracket$, the ideal $\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}^{r}_{\geq 0}\rangle\subset R$ is generated by $\{f_{\bm{n}}\mid\bm{n}\leq\bm{\ell}\}$ by Lemma 2.1. ∎
Note that when ${\mathrm{char}}R>0$, we can reduce the situation to the case where $\deg_{n_{j}}c_{k}<{\mathrm{char}}R$ for all $j$ and $k$.
@@
for $n\in{\mathbb{Z}}_{\geq 0}$. Since degree of the coefficients $1,1,1,1$ are all 0, we have $\ell=4-1=3$. By Theorem 2.5, $\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $f_{0},\dots,f_{3}$.
@@ 3. The case of higher degree exponents
and $I=\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}\rangle$. We consider a finite generating set of II. Fix a sutable $\bm{k}_{0}\in{\mathbb{Z}}_{\geq 0}^{r}$ and let $J=\langle f_{\bm{n}}\mid\bm{n}\leq\bm{k}_{0}\rangle$. Then take a sufficient large $\bm{k}_{1}\in{\mathbb{Z}}_{\geq 0}^{r}$ with $\bm{k}_{0}<\bm{k}_{1}$, and compute the list